214  PART 31: DEVICE-INDEPENDENT FILE FORMAT  TEX82 §583

583. Device-independent file format. The most important output produced by a run of TEX is the “device independent” ( ) file that specifies where characters and rules are to appear on printed pages. The form of these files was designed by David R. Fuchs in 1979. Almost any reasonable typesetting device can be driven by a program that takes files as input, and dozens of such -to-whatever programs have been written. Thus, it is possible to print the output of TEX on many different kinds of equipment, using TEX as a device-independent “front end.”

A file is a stream of 8-bit bytes, which may be regarded as a series of commands in a machine-like language. The first byte of each command is the operation code, and this code is followed by zero or more bytes that provide parameters to the command. The parameters themselves may consist of several consecutive bytes; for example, the ‘set rule’ command has two parameters, each of which is four bytes long. Parameters are usually regarded as nonnegative integers; but four-byte-long parameters, and shorter parameters that denote distances, can be either positive or negative. Such parameters are given in two’s complement notation. For example, a two-byte-long distance parameter has a value between −2^{15} and 2^{15} − 1. As in files, numbers that occupy more than one byte position appear in BigEndian order.

A file consists of a “preamble,” followed by a sequence of one or more “pages,” followed by a “postamble.” The preamble is simply a pre command, with its parameters that define the dimensions used in the file; this must come first. Each “page” consists of a bop command, followed by any number of other commands that tell where characters are to be placed on a physical page, followed by an eop command. The pages appear in the order that TEX generated them. If we ignore nop commands and fnt def commands (which are allowed between any two commands in the file), each eop command is immediately followed by a bop command, or by a post command; in the latter case, there are no more pages in the file, and the remaining bytes form the postamble. Further details about the postamble will be explained later.

Some parameters in commands are “pointers.” These are four-byte quantities that give the location number of some other byte in the file; the first byte is number 0, then comes number 1, and so on. For example, one of the parameters of a bop command points to the previous bop; this makes it feasible to read the pages in backwards order, in case the results are being directed to a device that stacks its output face up. Suppose the preamble of a file occupies bytes 0 to 99. Now if the first page occupies bytes 100 to 999, say, and if the second page occupies bytes 1000 to 1999, then the bop that starts in byte 1000 points to 100 and the bop that starts in byte 2000 points to 1000. (The very first bop, i.e., the one starting in byte 100, has a pointer of −1.)

584. The format is intended to be both compact and easily interpreted by a machine. Compactness is achieved by making most of the information implicit instead of explicit. When a -reading program reads the commands for a page, it keeps track of several quantities: (a) The current font f is an integer; this value is changed only by fnt and fnt num commands. (b) The current position on the page is given by two numbers called the horizontal and vertical coordinates, h and v. Both coordinates are zero at the upper left corner of the page; moving to the right corresponds to increasing the horizontal coordinate, and moving down corresponds to increasing the vertical coordinate. Thus, the coordinates are essentially Cartesian, except that vertical directions are flipped; the Cartesian version of (h, v) would be (h, −v). (c) The current spacing amounts are given by four numbers w, x, y, and z, where w and x are used for horizontal spacing and where y and z are used for vertical spacing. (d) There is a stack containing (h, v, w, x, y, z) values; the commands push and pop are used to change the current level of operation. Note that the current font f is not pushed and popped; the stack contains only information about positioning.

The values of h, v, w, x, y, and z are signed integers having up to 32 bits, including the sign. Since they represent physical distances, there is a small unit of measurement such that increasing h by 1 means moving a certain tiny distance to the right. The actual unit of measurement is variable, as explained below; TEX sets things up so that its output is in sp units, i.e., scaled points, in agreement with all the scaled dimensions in TEX’s data structures.

§585  TEX82  PART 31: DEVICE-INDEPENDENT FILE FORMAT 215

585. Here is a list of all the commands that may appear in a file. Each command is specified by its symbolic name (e.g., bop), its opcode byte (e.g., 139), and its parameters (if any). The parameters are followed by a bracketed number telling how many bytes they occupy; for example, ‘p[4]’ means that parameter p is four bytes long. set char 0 0. Typeset character number 0 from font f such that the reference point of the character is

at (h, v). Then increase h by the width of that character. Note that a character may have zero or negative width, so one cannot be sure that h will advance after this command; but h usually does increase.

set char 1 through set char 127 (opcodes 1 to 127). Do the operations of set char 0 ; but use the character

whose number matches the opcode, instead of character 0. set1 128 c[1]. Same as set char 0 , except that character number c is typeset. TEX82 uses this command for

characters in the range 128 ≤ c < 256. set2 129 c[2]. Same as set1 , except that c is two bytes long, so it is in the range 0 ≤ c < 65536. TEX82

never uses this command, but it should come in handy for extensions of TEX that deal with oriental languages.

set3 130 c[3]. Same as set1 , except that c is three bytes long, so it can be as large as 2^{24} − 1. Not even the Chinese language has this many characters, but this command might prove useful in some yet unforeseen extension.

set4 131 c[4]. Same as set1 , except that c is four bytes long. Imagine that. set rule 132 a[4] b[4]. Typeset a solid black rectangle of height a and width b, with its bottom left corner at

(h, v). Then set h ← h + b. If either a ≤ 0 or b ≤ 0, nothing should be typeset. Note that if b < 0, the value of h will decrease even though nothing else happens. See below for details about how to typeset rules so that consistency with is guaranteed.

put1 133 c[1]. Typeset character number c from font f such that the reference point of the character is at (h, v). (The ‘put’ commands are exactly like the ‘set’ commands, except that they simply put out a character or a rule without moving the reference point afterwards.)

put2 134 c[2]. Same as set2 , except that h is not changed. put3 135 c[3]. Same as set3 , except that h is not changed. put4 136 c[4]. Same as set4 , except that h is not changed. put rule 137 a[4] b[4]. Same as set rule, except that h is not changed. nop 138. No operation, do nothing. Any number of nop’s may occur between commands, but a nop

cannot be inserted between a command and its parameters or between two parameters. bop 139 c_{0}[4] c_{1}[4] … c_{9}[4] p[4]. Beginning of a page: Set (h, v, w, x, y, z) ← (0, 0, 0, 0, 0, 0) and set the

stack empty. Set the current font f to an undefined value. The ten c_{i} parameters hold the values of … in TEX at the time was invoked for this page; they can be used to identify pages, if a user wants to print only part of a file. The parameter p points to the previous bop in the file; the first bop has p = −1.

eop 140. End of page: Print what you have read since the previous bop. At this point the stack should be empty. (The -reading programs that drive most output devices will have kept a buffer of the material that appears on the page that has just ended. This material is largely, but not entirely, in order by v coordinate and (for fixed v) by h coordinate; so it usually needs to be sorted into some order that is appropriate for the device in question.)

push 141. Push the current values of (h, v, w, x, y, z) onto the top of the stack; do not change any of these

values. Note that f is not pushed. pop 142. Pop the top six values off of the stack and assign them respectively to (h, v, w, x, y, z). The number

of pops should never exceed the number of pushes, since it would be highly embarrassing if the stack were empty at the time of a pop command.

right1 143 b[1]. Set h ← h+b, i.e., move right b units. The parameter is a signed number in two’s complement notation, −128 ≤ b < 128; if b < 0, the reference point moves left.

216  PART 31: DEVICE-INDEPENDENT FILE FORMAT  TEX82 §585

right2 144 b[2]. Same as right1 , except that b is a two-byte quantity in the range −32768 ≤ b < 32768. right3 145 b[3]. Same as right1 , except that b is a three-byte quantity in the range −2^{23} ≤ b < 2^{23}. right4 146 b[4]. Same as right1 , except that b is a four-byte quantity in the range −2^{31} ≤ b < 2^{31}. w0 147. Set h ← h+w; i.e., move right w units. With luck, this parameterless command will usually suffice,

because the same kind of motion will occur several times in succession; the following commands explain how w gets particular values.

w1 148 b[1]. Set w ← b and h ← h + b. The value of b is a signed quantity in two’s complement notation,

−128 ≤ b < 128. This command changes the current w spacing and moves right by b. w2 149 b[2]. Same as w1 , but b is two bytes long, −32768 ≤ b < 32768. w3 150 b[3]. Same as w1 , but b is three bytes long, −2^{23} ≤ b < 2^{23}. w4 151 b[4]. Same as w1 , but b is four bytes long, −2^{31} ≤ b < 2^{31}. x0 152. Set h ← h + x; i.e., move right x units. The ‘x’ commands are like the ‘w’ commands except that

they involve x instead of w. x1 153 b[1]. Set x ← b and h ← h + b. The value of b is a signed quantity in two’s complement notation,

−128 ≤ b < 128. This command changes the current x spacing and moves right by b. x2 154 b[2]. Same as x1 , but b is two bytes long, −32768 ≤ b < 32768. x3 155 b[3]. Same as x1 , but b is three bytes long, −2^{23} ≤ b < 2^{23}. x4 156 b[4]. Same as x1 , but b is four bytes long, −2^{31} ≤ b < 2^{31}. down1 157 a[1]. Set v ← v + a, i.e., move down a units. The parameter is a signed number in two’s

complement notation, −128 ≤ a < 128; if a < 0, the reference point moves up. down2 158 a[2]. Same as down1 , except that a is a two-byte quantity in the range −32768 ≤ a < 32768. down3 159 a[3]. Same as down1 , except that a is a three-byte quantity in the range −2^{23} ≤ a < 2^{23}. down4 160 a[4]. Same as down1 , except that a is a four-byte quantity in the range −2^{31} ≤ a < 2^{31}. y0 161. Set v ← v + y; i.e., move down y units. With luck, this parameterless command will usually suffice,

because the same kind of motion will occur several times in succession; the following commands explain how y gets particular values.

y1 162 a[1]. Set y ← a and v ← v + a. The value of a is a signed quantity in two’s complement notation,

−128 ≤ a < 128. This command changes the current y spacing and moves down by a. y2 163 a[2]. Same as y1 , but a is two bytes long, −32768 ≤ a < 32768. y3 164 a[3]. Same as y1 , but a is three bytes long, −2^{23} ≤ a < 2^{23}. y4 165 a[4]. Same as y1 , but a is four bytes long, −2^{31} ≤ a < 2^{31}. z0 166. Set v ← v + z; i.e., move down z units. The ‘z’ commands are like the ‘y’ commands except that

they involve z instead of y. z1 167 a[1]. Set z ← a and v ← v + a. The value of a is a signed quantity in two’s complement notation,

−128 ≤ a < 128. This command changes the current z spacing and moves down by a. z2 168 a[2]. Same as z1 , but a is two bytes long, −32768 ≤ a < 32768. z3 169 a[3]. Same as z1 , but a is three bytes long, −2^{23} ≤ a < 2^{23}. z4 170 a[4]. Same as z1 , but a is four bytes long, −2^{31} ≤ a < 2^{31}. fnt num 0 171. Set f ← 0. Font 0 must previously have been defined by a fnt def instruction, as explained

below. fnt num 1 through fnt num 63 (opcodes 172 to 234). Set f ← 1, …, f ← 63, respectively. fnt1 235 k[1]. Set f ← k. TEX82 uses this command for font numbers in the range 64 ≤ k < 256. fnt2 236 k[2]. Same as fnt1 , except that k is two bytes long, so it is in the range 0 ≤ k < 65536. TEX82

never generates this command, but large font numbers may prove useful for specifications of color or texture, or they may be used for special fonts that have fixed numbers in some external coding scheme. fnt3 237 k[3]. Same as fnt1 , except that k is three bytes long, so it can be as large as 2^{24} − 1. fnt4 238 k[4]. Same as fnt1 , except that k is four bytes long; this is for the really big font numbers (and for

the negative ones). xxx1 239 k[1] x[k]. This command is undefined in general; it functions as a (k + 2)-byte nop unless special

-reading programs are being used. TEX82 generates xxx1 when a short enough appears, setting k to the number of bytes being sent. It is recommended that x be a string having the form of a keyword followed by possible parameters relevant to that keyword.

xxx2 240 k[2] x[k]. Like xxx1 , but 0 ≤ k < 65536. xxx3 241 k[3] x[k]. Like xxx1 , but 0 ≤ k < 2^{24}. xxx4 242 k[4] x[k]. Like xxx1 , but k can be ridiculously large. TEX82 uses xxx4 when sending a string of

length 256 or more. fnt def1 243 k[1] c[4] s[4] d[4] a[1] l[1] n[a + l]. Define font k, where 0 ≤ k < 256; font definitions will be

explained shortly. fnt def2 244 k[2] c[4] s[4] d[4] a[1] l[1] n[a + l]. Define font k, where 0 ≤ k < 65536. fnt def3 245 k[3] c[4] s[4] d[4] a[1] l[1] n[a + l]. Define font k, where 0 ≤ k < 2^{24}. fnt def4 246 k[4] c[4] s[4] d[4] a[1] l[1] n[a + l]. Define font k, where −2^{31} ≤ k < 2^{31}. pre 247 i[1] num[4] den[4] mag [4] k[1] x[k]. Beginning of the preamble; this must come at the very beginning

of the file. Parameters i, num, den, mag , k, and x are explained below. post 248. Beginning of the postamble, see below. post post 249. Ending of the postamble, see below. Commands 250–255 are undefined at the present time.

586. define set char 0 = 0 { typeset character 0 and move right } define set1 = 128 { typeset a character and move right } define set rule = 132 { typeset a rule and move right } define put rule = 137 { typeset a rule } define nop = 138 { no operation } define bop = 139 { beginning of page } define eop = 140 { ending of page } define push = 141 { save the current positions } define pop = 142 { restore previous positions } define right1 = 143 { move right } define w0 = 147 { move right by w } define w1 = 148 { move right and set w } define x0 = 152 { move right by x } define x1 = 153 { move right and set x } define down1 = 157 { move down } define y0 = 161 { move down by y } define y1 = 162 { move down and set y } define z0 = 166 { move down by z } define z1 = 167 { move down and set z } define fnt num 0 = 171 { set current font to 0 } define fnt1 = 235 { set current font } define xxx1 = 239 { extension to primitives } define xxx4 = 242 { potentially long extension to primitives } define fnt def1 = 243 { define the meaning of a font number } define pre = 247 { preamble } define post = 248 { postamble beginning } define post post = 249 { postamble ending } 587. The preamble contains basic information about the file as a whole. As stated above, there are six parameters:

i[1] num[4] den[4] mag [4] k[1] x[k].

The i byte identifies format; currently this byte is always set to 2. (The value i = 3 is currently used for an extended format that allows a mixture of right-to-left and left-to-right typesetting. Some day we will set i = 4, when format makes another incompatible change—perhaps in the year 2048.)

The next two parameters, num and den, are positive integers that define the units of measurement; they are the numerator and denominator of a fraction by which all dimensions in the file could be multiplied in order to get lengths in units of 10^{−7} meters. Since 7227pt = 254cm, and since TEX works with scaled points where there are 2^{16} sp in a point, TEX sets num/den = (254·10^{5})/(7227·2^{16}) = 25400000/473628672.

The mag parameter is what TEX calls , i.e., 1000 times the desired magnification. The actual fraction by which dimensions are multiplied is therefore mag · num/1000den. Note that if a TEX source document does not call for any ‘ ’ dimensions, and if you change it only by specifying a different setting, the file that TEX creates will be completely unchanged except for the value of mag in the preamble and postamble. (Fancy -reading programs allow users to override the mag setting when a file is being printed.)

Finally, k and x allow the writer to include a comment, which is not interpreted further. The length of comment x is k, where 0 ≤ k < 256.

define id byte = 2 { identifies the kind of files described here }

588. Font definitions for a given font number k contain further parameters

c[4] s[4] d[4] a[1] l[1] n[a + l].

The four-byte value c is the check sum that TEX found in the file for this font; c should match the check sum of the font found by programs that read this file.

Parameter s contains a fixed-point scale factor that is applied to the character widths in font k; font dimensions in files and other font files are relative to this quantity, which is called the “at size” elsewhere in this documentation. The value of s is always positive and less than 2^{27}. It is given in the same units as the other dimensions, i.e., in sp when TEX82 has made the file. Parameter d is similar to s; it is the “design size,” and (like s) it is given in units. Thus, font k is to be used at mag · s/1000d times its normal size.

The remaining part of a font definition gives the external name of the font, which is an ASCII string of length a + l. The number a is the length of the “area” or directory, and l is the length of the font name itself; the standard local system font area is supposed to be used when a = 0. The n field contains the area in its first a bytes.

Font definitions must appear before the first use of a particular font number. Once font k is defined, it must not be defined again; however, we shall see below that font definitions appear in the postamble as well as in the pages, so in this sense each font number is defined exactly twice, if at all. Like nop commands, font definitions can appear before the first bop, or between an eop and a bop.

589. Sometimes it is desirable to make horizontal or vertical rules line up precisely with certain features in characters of a font. It is possible to guarantee the correct matching between output and the characters generated by by adhering to the following principles: (1) The characters should be positioned so that a bottom edge or left edge that is supposed to line up with the bottom or left edge of a rule appears at the reference point, i.e., in row 0 and column 0 of the raster. This ensures that the position of the rule will not be rounded differently when the pixel size is not a perfect multiple of the units of measurement in the file. (2) A typeset rule of height a > 0 and width b > 0 should be equivalent to a -generated character having black pixels in precisely those raster positions whose coordinates satisfy 0 ≤ x < αb and 0 ≤ y < αa, where α is the number of pixels per unit.

590. The last page in a file is followed by ‘post ’; this command introduces the postamble, which summarizes important facts that TEX has accumulated about the file, making it possible to print subsets of the data with reasonable efficiency. The postamble has the form

post p[4] num[4] den[4] mag [4] l[4] u[4] s[2] t[2] ⟨ font definitions ⟩ post post q[4] i[1] 223’s[≥4]

Here p is a pointer to the final bop in the file. The next three parameters, num, den, and mag , are duplicates of the quantities that appeared in the preamble.

Parameters l and u give respectively the height-plus-depth of the tallest page and the width of the widest page, in the same units as other dimensions of the file. These numbers might be used by a -reading program to position individual “pages” on large sheets of film or paper; however, the standard convention for output on normal size paper is to position each page so that the upper left-hand corner is exactly one inch from the left and the top. Experience has shown that it is unwise to design -to-printer software that attempts cleverly to center the output; a fixed position of the upper left corner is easiest for users to understand and to work with. Therefore l and u are often ignored.

Parameter s is the maximum stack depth (i.e., the largest excess of push commands over pop commands) needed to process this file. Then comes t, the total number of pages (bop commands) present.

The postamble continues with font definitions, which are any number of fnt def commands as described above, possibly interspersed with nop commands. Each font number that is used in the file must be defined exactly twice: Once before it is first selected by a fnt command, and once in the postamble.

591. The last part of the postamble, following the post post byte that signifies the end of the font definitions, contains q, a pointer to the post command that started the postamble. An identification byte, i, comes next; this currently equals 2, as in the preamble.

The i byte is followed by four or more bytes that are all equal to the decimal number 223 (i.e., ´ 337 in octal). TEX puts out four to seven of these trailing bytes, until the total length of the file is a multiple of four bytes, since this works out best on machines that pack four bytes per word; but any number of 223’s is allowed, as long as there are at least four of them. In effect, 223 is a sort of signature that is added at the very end.

This curious way to finish off a file makes it feasible for -reading programs to find the postamble first, on most computers, even though TEX wants to write the postamble last. Most operating systems permit random access to individual words or bytes of a file, so the reader can start at the end and skip backwards over the 223’s until finding the identification byte. Then it can back up four bytes, read q, and move to byte q of the file. This byte should, of course, contain the value 248 (post ); now the postamble can be read, so the reader can discover all the information needed for typesetting the pages. Note that it is also possible to skip through the file at reasonably high speed to locate a particular page, if that proves desirable. This saves a lot of time, since files used in production jobs tend to be large.

Unfortunately, however, standard Pascal does not include the ability to access a random position in a file, or even to determine the length of a file. Almost all systems nowadays provide the necessary capabilities, so format has been designed to work most efficiently with modern operating systems. But if files have to be processed under the restrictions of standard Pascal, one can simply read them from front to back, since the necessary header information is present in the preamble and in the font definitions. (The l and u and s and t parameters, which appear only in the postamble, are “frills” that are handy but not absolutely necessary.)